QUESTION 2 (TOPIC: OVERHEAD COSTING)



Given: Total budgeted overheads = Shs.240, 000 and the Production budget is given as follows:


Product A B


i) Units 20,000 10,000


ii) Labour hours 20,000 20,000


iii) Labour cost Shs.17,500 Shs.22,500


iv) Machine hours 45,000 15,000


v) Material cost Shs.15,000 Shs.25,000



Required;


The overhead absorption rate per unit of A and B using the following methods:


a) Unit method


b) Percentage on material cost.


c) Percentage on labour cost.


d) Percentage On prime cost.


e) Labour hour rate.


f) Machine hour rate.


1
Expert's answer
2021-11-05T15:12:03-0400

a) Unit method


overhead absorption rate= "\frac{Total budgeted overheads}{Units}"


Of A: ="\frac{Shs240,000}{20,000}=Shs". 12 per unit of B: ="\frac{Shs240,000}{10,000}=Shs". 24 per unit



b) Percentage on material cost.


overhead absorption rate= "\frac{Total budgeted overheads}{Material costs}"



Of A: ="\frac{Shs240,000}{15,000}\times100\%=1600\%" Of B: ="\frac{Shs240,000}{25,000}\times100\%=960\%"



c) Percentage on labour cost.


overhead absorption rate= "\frac{Total budgeted overheads}{Labourcosts}"



Of A: ="\frac{Shs240,000}{17,500}\times100\%=1371.43\%" Of B: ="\frac{Shs240,000}{22,500}\times100\%=1066.67\%"


d) Percentage On prime cost.


overhead absorption rate= "\frac{Total budgeted overheads}{Primecosts}"


Of A: ="\frac{Shs240,000}{32,500}\times100\%=738.46\%" Of B: ="\frac{Shs240,000}{47,500}\times100\%=505.26\%"


e) Labour hour rate.


overhead absorption rate= "\frac{Total budgeted overheads}{labour hours}"


Of A: ="\frac{Shs240,000}{20,000}=Shs". 12 per unit Of B: ="\frac{Shs240,000}{20,000}=Shs". 12 per unit


f) Machine hour rate.


overhead absorption rate= "\frac{Total budgeted overheads}{machine hours}"


Of A: ="\frac{Shs240,000}{45,000}=Shs". 5.33 per machine hour Of B ="\frac{Shs240,000}{15,000}=Shs". 16 per machine hour